# Quark-Gluon Plasma, Charmonia, Ultra-relativistic Heavy-Ion Collisions

### v2

@@ Abstract
We provide an update on our semi-classical transport approach for quarkonium production in high-energy heavy-ion collisions, focusing on $J/\psi$ and $\psi(2S)$ mesons in 5.02 TeV Pb-Pb collisions at the Large Hadron Collider (LHC) at both forward and mid-rapidity. In particular, we employ the most recent charm-production cross sections reported in $p$ collisions, which are pivotal for the magnitude of the regeneration contribution, and their modifications due to cold-nuclear-matter (CNM) effects. Multi-differential observables are calculated in terms of nuclear modification factors as a function of centrality, transverse momentum, and rapidity, including the contributions from feeddown from bottom-hadron decays. For our predictions for $\psi(2S)$ production, the mechanism of sequential regeneration relative to the more strongly bound $J/\psi$ meson plays an important role in interpreting recent ALICE data.
@@ 1 Introduction
The production of charmonia in ultra-relativistic heavy-ion collisions (URHICs) is an active area of research since four decades ago. The initially proposed $J/\psi$ suppression signature of quark-gluon plasma (QGP) formation Matsui:1986dk has developed into more comprehensive transport models that account for regeneration mechanisms as dictated by the principle of detailed balance which ensures that the abundances of charmonia approach their pertinent equilibrium limits, see, e.g., Refs. Rapp:2008tf; Braun-Munzinger:2009dzl; Liu:2015izf; Andronic:2024oxz for reviews. Abundant charm production at the Large Hadron Collider (LHC), with around 100 charm-anticharm quark pairs in a central Pb-Pb collision at a center-of-mass energy of 5.02 TeV per nucleon pair, has led to predictions of a substantial amount of regenerated charmonia which have been confirmed by experiment ALICE:2016flj; Bai:2020svs; ALICE:2023gco; ALICE:2022jeh; ALICE:2019lga; ALICE:2019nrq. This signature features an approximately constant (or even rising) $J/\psi$ yield with collision centrality in terms of the nuclear modification factor, $R_{\rm AA}$, and a concentration of the regeneration yield at relatively low momenta Zhao:2011cv; Song:2011xi; Zhou:2014kka; Ferreiro:2014bia. In addition, the regenerated charmonia exhibit an appreciable elliptic flow inherited from the recombining charm and anti-charm quarks that have been dragged along with the expanding fireball He:2021zej. However, significant model uncertainties remain, most notably in the underlying assumptions about the $J/\psi$ dissociation temperature, which controls the onset of regeneration in the cooling fireball, and in the input charm cross section, which determines the equilibrium limit of the charmonia and thus controls the magnitude of the regeneration. For example, the Statistical Hadronization Model (SHM) offers a complementary perspective, where all charmonium states are produced via statistical hadronization at a fixed common temperature corresponding the pseudo-critical temperature of the chiral cross-over transition, $T_{\rm pc}\simeq 160$ MeV Andronic:2021erx (see also Ref. Capellino:2023cxe), while most transport models are based on a hierarchy of dissociation temperatures that is correlated with the charmonium binding energies. Both transport and SHM models are quite sensitive to the amount of charm-anti-charm quark pairs in the fireball. Fortunately, the experimental knowledge about the total charm cross section in proton-proton ($p$) collisions has much advanced in recent years. With higher precision and an improved assessment of the contribution of charm baryons a noticeable increase in the value of the cross section has emerged ALICE:2021dhb.
@@
This paper is organized as follows: In Sec. 2, we briefly recall our calculations of charmonium reaction rates and how they figure in the kinetic-rate equation within a schematic fireball for Pb-Pb collisions at the LHC; we also pinpoint updates specific to this work, e.g., in-medium charmonium binding energies, the total charm cross section with corrections from nuclear shadowing and bottom-decay feeddown in the nuclear modification factors. In Sec. 3 we discuss the time dependence of charmonia kinetics in central and peripheral Pb-Pb collisions at the LHC, as well as the centrality dependence of their inclusive yields with comparisons to data. In Sec. 4, we evaluate the $p_{\rm T}$ dependence of $J/\psi$ and $\psi(2S)$ production in Pb-Pb (5.02 TeV) collisions, based on fits to $p$ spectra. This analysis encompasses $p_{\rm T}$ spectra, pertinent nuclear modification factors across different centralities, centrality dependent yields within different momentum bins, and the average $p_{\rm T}$ its square, with comparisons to data as available. We summarize and conclude in Sec. 5.
@@ 2.4 Initial Conditions and Medium Evolution
Direct-production cross section ratio in $p$ collisions
@@
The cross section for $c\bar{c}$ pair production in $p$ collisions is a key input for the equilibrium limits and thus controls the amount of regeneration. It is usually applied for a specific rapidity interval as $N_{c\bar{c}}=\mathrm{d}\sigma_{c\bar{c}}/\mathrm{d}yN_{\rm coll}$, where $N_{\rm coll}$ denotes the number of binary nucleon-nucleon collisions at a given collision energy and centrality (estimated from the optical Glauber model Miller:2007ri; for the partonic production processes under consideration here, which are mostly driven by gluon fusion, no distinction is made for the proton-neutron and neutron-neutron collisions). Similarly, we obtain the initial number of charmonium states, which are required to determine the initial condition for the rate equation. In Table 1 we summarize the charm/onium cross sections for $p$ collisions at 5.02 TeV and the ratio of $\psi(2S)$ to $J/\psi$ (we note that our values for the charm cross section, taken from Ref. ALICE:2021dhb, are well in line with the most recent assessment in Ref. ALICE:2023sgl). The open-charm cross section at forward rapidity has been deduced from its counterpart at midrapidity through the rapidity dependence outlined in Ref. Bierlich:2023ewv. This extrapolation yields a value $0.72\pm 0.07$, where we incorporate an uncertainty of $\pm 10\%$.
@@
In our comparisons to experimental results for Pb-Pb collisions discussed below, we incorporate the experimental uncertainties of the charm-anticharm cross section $\sigma_{c\bar{c}}$, denoted by $\Delta\sigma_{c\bar{c}}$, along with the shadowing factor $S$ and its uncertainty, $\Delta S$. The uncertainty of the $p$ cross section and its shadowing is combined into an effective cross section, $\tilde{\sigma}_{c\bar{c}}$, as follows:
@@
Figure 5: Fraction of bottom-hadron decay feeddown to inclusive $J/\psi$ production as measured by ALICE, ATLAS and CMS in $p$ collisions at the LHC Andronic:2015wma; our functional fit is shown by the solid line.
Data for inclusive $J/\psi$ production also include feeddown contributions from final-state decays. For the “prompt" feeddown from electromagnetic and strong decays of excited states, we account for 8% and 25% from $\psi(2S)$ and $\chi_{c}$ mesons, respectively Andronic:2015wma. Additionally, there is also a “non-prompt" fraction from bottom decay feeddown (unless explicitly removed experimentally). We estimate this fraction from available data Andronic:2015wma in $p$ collisions at the LHC, cf. Figure 5. It amounts to around 5% at $p_{\rm T}$=0 and increases to approximately 50% at $p_{\rm T}=20\,{\rm GeV}/c$. We fit the data using the empirical parameterization
@@
Right after the near-instantaneous formation of a $c\bar{c}$ pair which would develop into a $\psi$ in a $p$ collision, we incorporate initial formation time effects to approximate the (quantum) expansion dynamics of evolving into a fully formed bound state. Contrary to expectations derived from a classical cross section perspective where the transverse size (i.e., cross sectional area) would grow quadratically with time, we utilize a scaling that is linear with time Farrar:1988me and scale down the reaction rates by a factor of $\tau/\tau_{\rm form}$ for $\tau\leq\tau_{\rm form}$. The quantum formation times are estimated based on energy uncertainties associated with the splitting in binding energies, i.e., $\tau_{\mathrm{form}}(J/\psi,\psi(2S),\chi_{c})=1,2,2\,\text{fm}/c$, respectively.
@@ 3 Time and Centrality Dependence of Charmonium Yields at the LHC
where $N_{\psi}^{pp}$ denotes the inclusive $\psi$ yield in $p$ collisions at the same collision energy. The number of nucleon participants, $N_{\rm part}$, is estimated from a Glauber model for a given impact parameter, $b$ Miller:2007ri, and serves as a measure of the centrality of the nuclear collision. Unless otherwise stated, the denominator of the $R_{\rm AA}$ will include both prompt and non-prompt feeddown contributions and utilize the central values for the input cross sections at given rapidity, and the reaction rates are for quasifree dissociation.
@@ 3.2 Centrality dependence
Let us also briefly come back to the effect of the updated binding energies as compared to our previous calculations, as mentioned in the first paragraph of Sec. $2.$ The impact of the somewhat larger binding energies (most notably around $T\simeq$220 MeV) on regeneration is negligible, as one finds an almost complete compensation between the reduced rates and the larger equilibrium limits. The smaller rates do, however, affect the primordial suppression. The maximal effect for central Pb-Pb collisions amounts to an increase of the primordial contribution by about 0.03 units in terms of the $R_{\rm AA}$.
@@ 4 Transverse-Momentum Spectra
We solve for the primordial part by employing the Boltzmann equation without the gain term, with initial conditions obtained from $p$ collisions as specified in Sec. 4.1). The yield from regeneration then follows from the difference of the homogeneous solutions and the full rate equation, and we assume its $p_{\rm T}$ dependence to be given by a thermal-blastwave expression for an average regeneration temperature based on our expansion model (Sec. 4.2). This approximation has, of course, its limitations, and we will discuss evidence for that in systematic comparisons to experimental data which will be carried out in Sec. 4.3.
@@ 4.1 Initial $p_{\rm T}$ spectra and their suppression in heavy-ion collisions
Figure 11: Normalized $p_{\rm T}$ spectra fitted to ALICE data of inclusive $J/\psi$ production in $p$ collisions at mid-rapidity ALICE:2019pid; ATLAS:2017prf; CMS:2017exb (left panel), forward rapidity ALICE:2021qlw (middle panel), and inclusive $\psi(2S)$ at forward rapidity ALICE:2021qlw.
To construct the initial conditions of charmonia, we first perform fits to their $p_{\rm T}$ spectra in $p$ (5.02 TeV) collisions from the ALICE ALICE:2019pid; ALICE:2021qlw, ATLAS ATLAS:2017prf and CMS CMS:2017exb collaborations based on the ansatz
@@ 4.2 Transverse-momentum spectra from regeneration
Figure 12: Transverse-momentum spectra (upper panels) and pertinent nuclear-modification factors (lower panels) of inclusive $J/\psi$ production in central (left panels) and semi-central (right panels) Pb-Pb (5.02 TeV) collisions at mid-rapidity, compared to ALICE data Bai:2020svs; ALICE:2023gco. The spectra in $p$ collisions, scaled by the pertinent binary-collision number, $N_{\rm coll}(b)$, are shown as dashed lines in the upper panels. The bands and the colors of the data have the same meaning as in Figure 8.
@@ 4.3 Comparison to experimental $p_{\rm T}$ spectra
The $p_{\rm T}$ dependent ratio of $\psi(2S)$ to $J/\psi$ yields at forward rapidity, as depicted in the left panel of Figure 17, also aligns reasonably well with the ALICE data ALICE:2022jeh. The increasing trend with $p_{\rm T}$ from our fits to $p$ data tends to overestimate the data for the ratio at the highest $p_{\rm T}$. This discrepancy also migrates into the AA result (which may be partly due to the underestimation of the $J/\psi$ production at intermediate $p_{\rm T}$). Indeed, the agreement is better when dividing out the $p$ reference spectra in the $R_{\rm AA}$ double ratio, shown in the right panel of Figure 17. Here, the slight deficit in the $J/\psi$ yield at intermediate $p_{\rm T}$ shows up as a mild maximum structure which is not observed in the data at this point. In principle, such a maximum could be another signature of sequential regeneration in central collisions, when the open-charm spectra are close to thermal equilibrium.
@@
Figure 19: $\langle p_{\rm T}\rangle$ (upper left) and $\langle p_{\rm T}^{2}\rangle$ (upper right) of $J/\psi$ in 5.02 TeV $p$ and Pb-Pb collisions at mid-rapidity. Their ratios relative to the $p$ results are shown in the lower panels; ALICE data are from Refs. ALICE:2022jeh; ALICE:2019nrq.
Finally, we attend to the centrality dependence of the average charmonium momentum, $\langle p_{\rm T}\rangle$, and its square, $\langle p_{\rm T}^{2}\rangle$. We calculate a “small" nuclear-modification factor" defined in Ref. Zhou:2014kka, normalized to the values in $p$ and Pb-Pb collisions,

### v3

@@ Abstract
We provide an update on our semi-classical transport approach for quarkonium production in high-energy heavy-ion collisions, focusing on $J/\psi$ and $\psi(2S)$ mesons in 5.02 TeV Pb-Pb collisions at the Large Hadron Collider (LHC) at both forward and mid-rapidity. In particular, we employ the most recent charm-production cross sections reported in pp collisions, which are pivotal for the magnitude of the regeneration contribution, and their modifications due to cold-nuclear-matter (CNM) effects. Multi-differential observables are calculated in terms of nuclear modification factors as a function of centrality, transverse momentum, and rapidity, including the contributions from feeddown from bottom-hadron decays. For our predictions for $\psi(2S)$ production, the mechanism of sequential regeneration relative to the more strongly bound $J/\psi$ meson plays an important role in interpreting recent ALICE data.
@@ 1 Introduction
The production of charmonia in ultra-relativistic heavy-ion collisions (URHICs) is an active area of research since four decades ago. The initially proposed $J/\psi$ suppression signature of quark-gluon plasma (QGP) formation Matsui:1986dk has developed into more comprehensive transport models that account for regeneration mechanisms as dictated by the principle of detailed balance which ensures that the abundances of charmonia approach their pertinent equilibrium limits, see, e.g., Refs. Rapp:2008tf; Braun-Munzinger:2009dzl; Liu:2015izf; Andronic:2024oxz for reviews. Abundant charm production at the Large Hadron Collider (LHC), with around 100 charm-anticharm quark pairs in a central Pb-Pb collision at a center-of-mass energy of 5.02 TeV per nucleon pair, has led to predictions of a substantial amount of regenerated charmonia which have been confirmed by experiment ALICE:2016flj; Bai:2020svs; ALICE:2023gco; ALICE:2022jeh; ALICE:2019lga; ALICE:2019nrq. This signature features an approximately constant (or even rising) $J/\psi$ yield with collision centrality in terms of the nuclear modification factor, $R_{\rm AA}$, and a concentration of the regeneration yield at relatively low momenta Zhao:2011cv; Song:2011xi; Zhou:2014kka; Ferreiro:2014bia. In addition, the regenerated charmonia exhibit an appreciable elliptic flow inherited from the recombining charm and anti-charm quarks that have been dragged along with the expanding fireball He:2021zej. However, significant model uncertainties remain, most notably in the underlying assumptions about the $J/\psi$ dissociation temperature, which controls the onset of regeneration in the cooling fireball, and in the input charm cross section, which determines the equilibrium limit of the charmonia and thus controls the magnitude of the regeneration. For example, the Statistical Hadronization Model (SHM) offers a complementary perspective, where all charmonium states are produced via statistical hadronization at a fixed common temperature corresponding the pseudo-critical temperature of the chiral cross-over transition, $T_{\rm pc}\simeq 160$ MeV Andronic:2021erx (see also Ref. Capellino:2023cxe), while most transport models are based on a hierarchy of dissociation temperatures that is correlated with the charmonium binding energies. Both transport and SHM models are quite sensitive to the amount of charm-anti-charm quark pairs in the fireball. Fortunately, the experimental knowledge about the total charm cross section in proton-proton (pp) collisions has much advanced in recent years. With higher precision and an improved assessment of the contribution of charm baryons a noticeable increase in the value of the cross section has emerged ALICE:2021dhb.
@@
This paper is organized as follows: In Sec. 2, we briefly recall our calculations of charmonium reaction rates and how they figure in the kinetic-rate equation within a schematic fireball for Pb-Pb collisions at the LHC; we also pinpoint updates specific to this work, e.g., in-medium charmonium binding energies, the total charm cross section with corrections from nuclear shadowing and bottom-decay feeddown in the nuclear modification factors. In Sec. 3 we discuss the time dependence of charmonia kinetics in central and peripheral Pb-Pb collisions at the LHC, as well as the centrality dependence of their inclusive yields with comparisons to data. In Sec. 4, we evaluate the $p_{\rm T}$ dependence of $J/\psi$ and $\psi(2S)$ production in Pb-Pb (5.02 TeV) collisions, based on fits to pp spectra. This analysis encompasses $p_{\rm T}$ spectra, pertinent nuclear modification factors across different centralities, centrality dependent yields within different momentum bins, and the average $p_{\rm T}$ its square, with comparisons to data as available. We summarize and conclude in Sec. 5.
@@ 2.4 Initial Conditions and Medium Evolution
Direct-production cross section ratio in pp collisions
@@
The cross section for $c\bar{c}$ pair production in pp collisions is a key input for the equilibrium limits and thus controls the amount of regeneration. It is usually applied for a specific rapidity interval as $N_{c\bar{c}}=\mathrm{d}\sigma_{c\bar{c}}/\mathrm{d}yN_{\rm coll}$, where $N_{\rm coll}$ denotes the number of binary nucleon-nucleon collisions at a given collision energy and centrality (estimated from the optical Glauber model Miller:2007ri; for the partonic production processes under consideration here, which are mostly driven by gluon fusion, no distinction is made for the proton-neutron and neutron-neutron collisions). Similarly, we obtain the initial number of charmonium states, which are required to determine the initial condition for the rate equation. In Table 1 we summarize the charm/onium cross sections for pp collisions at 5.02 TeV and the ratio of $\psi(2S)$ to $J/\psi$ (we note that our values for the charm cross section, taken from Ref. ALICE:2021dhb, are well in line with the most recent assessment in Ref. ALICE:2023sgl). The open-charm cross section at forward rapidity has been deduced from its counterpart at midrapidity through the rapidity dependence outlined in Ref. Bierlich:2023ewv. This extrapolation yields a value $0.72\pm 0.07$, where we incorporate an uncertainty of $\pm 10\%$.
@@
In our comparisons to experimental results for Pb-Pb collisions discussed below, we incorporate the experimental uncertainties of the charm-anticharm cross section $\sigma_{c\bar{c}}$, denoted by $\Delta\sigma_{c\bar{c}}$, along with the shadowing factor $S$ and its uncertainty, $\Delta S$. The uncertainty of the pp cross section and its shadowing is combined into an effective cross section, $\tilde{\sigma}_{c\bar{c}}$, as follows:
@@
Figure 5: Fraction of bottom-hadron decay feeddown to inclusive $J/\psi$ production as measured by ALICE, ATLAS and CMS in pp collisions at the LHC Andronic:2015wma; our functional fit is shown by the solid line.
Data for inclusive $J/\psi$ production also include feeddown contributions from final-state decays. For the “prompt" feeddown from electromagnetic and strong decays of excited states, we account for 8% and 25% from $\psi(2S)$ and $\chi_{c}$ mesons, respectively Andronic:2015wma. Additionally, there is also a “non-prompt" fraction from bottom decay feeddown (unless explicitly removed experimentally). We estimate this fraction from available data Andronic:2015wma in pp collisions at the LHC, cf. Figure 5. It amounts to around 5% at $p_{\rm T}$=0 and increases to approximately 50% at $p_{\rm T}=20\,{\rm GeV}/c$. We fit the data using the empirical parameterization
@@
Right after the near-instantaneous formation of a $c\bar{c}$ pair which would develop into a $\psi$ in a pp collision, we incorporate initial formation time effects to approximate the (quantum) expansion dynamics of evolving into a fully formed bound state. Contrary to expectations derived from a classical cross section perspective where the transverse size (i.e., cross sectional area) would grow quadratically with time, we utilize a scaling that is linear with time Farrar:1988me and scale down the reaction rates by a factor of $\tau/\tau_{\rm form}$ for $\tau\leq\tau_{\rm form}$. The quantum formation times are estimated based on energy uncertainties associated with the splitting in binding energies, i.e., $\tau_{\mathrm{form}}(J/\psi,\psi(2S),\chi_{c})=1,2,2\,\text{fm}/c$, respectively.
@@ 3 Time and Centrality Dependence of Charmonium Yields at the LHC
where $N_{\psi}^{pp}$ denotes the inclusive $\psi$ yield in pp collisions at the same collision energy. The number of nucleon participants, $N_{\rm part}$, is estimated from a Glauber model for a given impact parameter, $b$ Miller:2007ri, and serves as a measure of the centrality of the nuclear collision. Unless otherwise stated, the denominator of the $R_{\rm AA}$ will include both prompt and non-prompt feeddown contributions and utilize the central values for the input cross sections at given rapidity, and the reaction rates are for quasifree dissociation.
@@ 3.2 Centrality dependence
Let us also briefly come back to the effect of the updated binding energies as compared to our previous calculations, as mentioned in the first paragraph of Sec. 2.2. The impact of the somewhat larger binding energies (most notably around $T\simeq$220 MeV) on regeneration is negligible, as one finds an almost complete compensation between the reduced rates and the larger equilibrium limits. The smaller rates do, however, affect the primordial suppression. The maximal effect for central Pb-Pb collisions amounts to an increase of the primordial contribution by about 0.03 units in terms of the $R_{\rm AA}$.
@@ 4 Transverse-Momentum Spectra
We solve for the primordial part by employing the Boltzmann equation without the gain term, with initial conditions obtained from pp collisions as specified in Sec. 4.1). The yield from regeneration then follows from the difference of the homogeneous solutions and the full rate equation, and we assume its $p_{\rm T}$ dependence to be given by a thermal-blastwave expression for an average regeneration temperature based on our expansion model (Sec. 4.2). This approximation has, of course, its limitations, and we will discuss evidence for that in systematic comparisons to experimental data which will be carried out in Sec. 4.3.
@@ 4.1 Initial $p_{\rm T}$ spectra and their suppression in heavy-ion collisions
Figure 11: Normalized $p_{\rm T}$ spectra fitted to ALICE data of inclusive $J/\psi$ production in pp collisions at mid-rapidity ALICE:2019pid; ATLAS:2017prf; CMS:2017exb (left panel), forward rapidity ALICE:2021qlw (middle panel), and inclusive $\psi(2S)$ at forward rapidity ALICE:2021qlw.
To construct the initial conditions of charmonia, we first perform fits to their $p_{\rm T}$ spectra in pp (5.02 TeV) collisions from the ALICE ALICE:2019pid; ALICE:2021qlw, ATLAS ATLAS:2017prf and CMS CMS:2017exb collaborations based on the ansatz
@@ 4.2 Transverse-momentum spectra from regeneration
Figure 12: Transverse-momentum spectra (upper panels) and pertinent nuclear-modification factors (lower panels) of inclusive $J/\psi$ production in central (left panels) and semi-central (right panels) Pb-Pb (5.02 TeV) collisions at mid-rapidity, compared to ALICE data Bai:2020svs; ALICE:2023gco. The spectra in pp collisions, scaled by the pertinent binary-collision number, $N_{\rm coll}(b)$, are shown as dashed lines in the upper panels. The bands and the colors of the data have the same meaning as in Figure 8.
@@ 4.3 Comparison to experimental $p_{\rm T}$ spectra
The $p_{\rm T}$ dependent ratio of $\psi(2S)$ to $J/\psi$ yields at forward rapidity, as depicted in the left panel of Figure 17, also aligns reasonably well with the ALICE data ALICE:2022jeh. The increasing trend with $p_{\rm T}$ from our fits to pp data tends to overestimate the data for the ratio at the highest $p_{\rm T}$. This discrepancy also migrates into the AA result (which may be partly due to the underestimation of the $J/\psi$ production at intermediate $p_{\rm T}$). Indeed, the agreement is better when dividing out the pp reference spectra in the $R_{\rm AA}$ double ratio, shown in the right panel of Figure 17. Here, the slight deficit in the $J/\psi$ yield at intermediate $p_{\rm T}$ shows up as a mild maximum structure which is not observed in the data at this point. In principle, such a maximum could be another signature of sequential regeneration in central collisions, when the open-charm spectra are close to thermal equilibrium.
@@
Figure 19: $\langle p_{\rm T}\rangle$ (upper left) and $\langle p_{\rm T}^{2}\rangle$ (upper right) of $J/\psi$ in 5.02 TeV pp and Pb-Pb collisions at mid-rapidity. Their ratios relative to the pp results are shown in the lower panels; ALICE data are from Refs. ALICE:2022jeh; ALICE:2019nrq.
Finally, we attend to the centrality dependence of the average charmonium momentum, $\langle p_{\rm T}\rangle$, and its square, $\langle p_{\rm T}^{2}\rangle$. We calculate a “small" nuclear-modification factor" defined in Ref. Zhou:2014kka, normalized to the values in pp and Pb-Pb collisions,